D Question 6 10.5 pts Charge 1 is positive and is located at the origin and charge 2 is negative. The electric force on charge 1 due to charge 2 is 319.5 degrees measured counter-clockwise from the +x axis. What is the direction of charge 2 due to charge 1 measured counter-clockwise from the +x axis, in degrees? Nothing related to the charges, changes... .they are still the same charges in the same positions. D Question 7 10.5 pts
Answer #1

From Newton's third law,

For every action there is equal and opposite reaction.

Force on charge 2 due to charge 1 = -Force on charge 1 due to charge 1

here -ve sign represents that direction of force will be opposite.

Now given that

Direction of Force on charge 1 due to charge 2 = 319.5 deg CCW from +x-axis

Direction of Force on charge 2 due to charge 1 = \theta deg CCW from +x-axis

Now these forces are in opposite directions, So

319.5 - \theta = 180 deg

\theta = 319.5 - 180

\theta = 139.5 deg Counter-clockwise from +x axis

Please Upvote.